OpenStudy (anonymous):

Where does the graph of y = 2x2 + 7x + 3 cross the x-axis

OpenStudy (anonymous):

you meant y= 2x^2+7x+3. substitute x=0, y=2(0)^2+7(0)+3 when x=0, y=3. you can find y values by substituting x=...,-2,-1,1,2,3...